### v1

@@ 3.4 Evaluation of OMR Models
Having discussed the benefits of having a common notation format and the importance of producing a structured output, we now delve into the specifics of evaluation of OMR systems.
Unfortunately, most metrics used by the OMR community in modern day are fairly domain-specific and rarely take into account music semantics, a direct consequence of not targeting a musically-meaningful output format Mengarelli et al (2019). Authors that address OMR as an object detection problem report object detection metrics, whereas authors that address OMR as an image-to-sequence task rely on sequence edit distances. Consequently, there is no realistic way of comparing them since the semantics of the mistakes of the system are lost and the underlying notations are different. The few authors that do report musically-aware metrics restrict themselves to per-note pitch and/or time accuracy Baró et al (2019); Huang et al (2019).
Moreover, end users for OMR systems are not concerned so much about the absolute number of mistakes, but rather how costly or realistic it is to fix them, something that is hard to convey with metrics that treat all errors equally. A simple example of why this distinction is relevant is the following: imagine there are two available OMR systems, one that is unreliable providing correct note pitches but is otherwise very robust and another one which is very good at determining music semantics but tends to miss structural primitives. Instead of devoting efforts improving both systems independently, the outputs of both models can be combined and weighted according to their strengths to produce better results at essentially zero additional cost. This is something that has already been done for other similar tasks such as Handwritten Text Recognition Bertolami and Bunke (2008).
In this direction, there is also debate regarding the specific aspects of music that should be measured. Hajić jr. Hajič jr. (2018) argues that the best metric is one that summarises the entire notation system as one single value, reminiscent of the Symbol Error Rate (SER) metric for strings. We believe that, while such a metric is a good indicator of overall performance (and is definitely useful for benchmarking), it must not be considered on its own, as definite conclusions about the underlying models are very hard to make. Our belief is that there should be an array of metrics that are oriented at different aspects of the score that should be logged together, increasing the coverage of the system.
@@ 3.5 The Ideal Scope for Recognition
At the current day there is no pre-defined scope to tackle the problem of OMR. Therefore researchers can choose whether to implement recognition at measure Baró et al (2020, 2022), line Calvo-Zaragoza and Rizo (2018a); Baró et al (2019); Wen and Zhu (2022) or page level Pacha et al (2018); Ríos-Vila et al (2022a); Zhang et al (2023). While there are practical advantages on each perspective, there is an arguing point that should be brought to discussion given the current situation of OMR and Computer Vision as a whole.
Current state-of-the-art object detection and recognition models struggle with both small Yu et al (2020); Pacha (2021); Baró et al (2022) and scale-variant symbols Zhang et al (2018). Music scores are quintessential examples of both phenomena: large symbols (slurs, ties, staves) coexist with very small ones (dots, fingering numbers), and writers draw symbols at their preferred relative scale without affecting their meaning – something that is particularly true for noteheads and stems – while also writing semantically-distinct symbols that change based on their size – grace notes. Moreover, the number of objects present in the input image also affects the recognition performance, a recurrent matter for a significant amount of pieces.
Since addressing visual recognition tasks at smaller scopes helps to obtain better results, we believe that OMR should first be approached at measure level, since this is the smallest semantic unit of music. The first immediate consequence is that layout analysis of music scores becomes a lot more important, as all successive recognition steps depend on finding the measures reliably on the page. Nevertheless, we argue that it is objectively a simpler problem than fully-fledged recognition: staves and bar lines are very prominent, very ubiquitous and can easily be identified heuristically Fujinaga (2004); Cardoso et al (2009); Fornés et al (2013); Bui et al (2014); Calvo-Zaragoza et al (2017a); Egozy and Clester (2022), as opposed to many of the other musical symbols. Therefore, by composing small semantic units and propagating shared semantics (key, time signature, clef) whenever needed, a full notation reconstruction can be obtained quite simply, as measures are mostly independent.
Tangentially, this idea also helps in the recognition by providing higher dpi images which are more informative, as there is a substantial part of the score that is usually plain white space. Some tricks have been developed to try to minimise these gaps, such as removing all horizontal space between primitives Zhang et al (2023). This is not something particularly useful with scores with noisy backgrounds (e.g. scans from old documents with signs of paper degradation, a recurrent matter within old music scores), but it has been proven to improve results substantially.
@@ 4 The Music Tree Notation Format
  • enables comparison of diverse OMR methods on equal grounds,
@@
  • simplifies conversion to a final structured format and
@@ 4.1 Rationale
The basis for our approach to OMR evaluation and (optionally) end-to-end inference is the Music Tree Notation format (MTN). We believe the final goal of structured OMR is the reconstruction of the score at the visual domain. The core idea of this format is therefore to build a notation that exclusively models relationships between graphical symbols and defers inference of music semantics until a later stage. Only those high-level music concepts that are strictly required to reconstruct the score unambiguously are kept if and only if there is a direct graphical cue that allows straightforward inference.
Evaluation should be possible regardless of the scope of recognition of the evaluated system – page, line or measure. Nevertheless, a minimal semantic unit is still required in order to reconstruct music semantics afterwards and facilitate alignment of predictions and their groundtruth. Exploiting the assumption evaluation is performed on the visual domain and therefore application of semantic-altering objects is irrelevant, we propose using the measure as the minimal unit of music on which to evaluate.
MTN should act as a final target for OMR, through which to convert the score into any other high-level notation system. This is in order to facilitate OMR systems from research being deployed into real use-cases as well as enabling any source of structured data to be used in OMR applications easily. Therefore, MTN draws heavy inspiration from both MEI and MusicXML, but “normalises” their structure in such a way that every score can only be expressed in one way. This also allows implementation many of the already-studied evaluation requirement ideas from other authors, particularly Hajič jr. Hajič jr. et al (2016) and Byrd Byrd and Simonsen (2015).
Finally, music notation is very rich and tends to bend its own rules quite heavily. The goal for this notation and evaluation system is to accommodate a most significant subset of CWMN, but covering every single exception proves to be a daunting task. Byrd Byrd and Simonsen (2015) presents some specific cases such as a spanning slur with seven inflection points that is rather complicated to model in the logical domain. Therefore, our rationale is that support for these should not be prioritary. We believe it is better to have some solid tools on a core subset of music rather than trying to cover every possible case with highly complex and fragile ones as this makes adoption of the format much more difficult.
@@ 4.2 Description of the Format
A graphical representation of a simple measure engraved in MTN can be seen in Figure 2. The core element of this format is the Musical Primitive, a concept that is quite widespread in the OMR literature Baró et al (2016, 2020); Calvo-Zaragoza and Rizo (2018a); Tuggener et al (2018) and can be defined as any of the independent structural elements that may or may not be combined together to form a semantic unit in the music score. The set of musical primitives includes all graphical elements in a score that are self-contained and require no other symbols to convey meaning (this includes rests, clefs or time signatures), the set of graphical elements that compose notes (noteheads, stems, flags, dots, etc.) and other miscellaneous elements such as numbers for compound time signatures. Every primitive is given a unique work-level identifier.
These primitives associate together to form more abstract constructs. This is modelled in MTN using a tree-like structure of higher-order elements resembling an Abstract Syntax Tree (AST), which defines the set of dependencies among objects in the score. This idea emulates parsing the contents of the score using a grammar, enabling the bulk of tools and research on parsers, parser generators and AST analysis and processing to be used in the context of music. Furthermore, it is a structure that can be modelled very easily using widespread exchange formats such as XML or JSON. The structure of the format may be found in Table 2.
Table 2: Definition of elements within MTN and the overall structure of the format. Each of the lowest levels of the notation contain the various final tokens. For more information, refer to the implementation of the format.
@@
Attributes
Any semantic change the score goes through.
Staff
Group of changes occuring on the same staff.
Clef
A change of clef in the score.
TimeSig
A change of time signature in the score.
Key
A change of key in the score.
Barline
A measure separator and its attached modifiers.
Direction
A single playback directive.
Note Group
Group of notes tied together by at leas one beam.
Chord
Set of notes in unison attached to the same stem.
Stem
Stem direction information and beams or flags attached to it.
Note
A notehead and its attached modifiers.
Rest
Rest token and its attached modifiers.
There are, however, some elements in music that break the tree-like structure assumption. These are elements that connect multiple notes together outside their local note group structure: slurs, ties, parentheses and tuplets, among others. Both MEI and MusicXML acknowledge this limitation and circumvent it through the use of identifiers. MTN is no different: it provides a unique starting and ending token for each side of the object and gives both ends the same identifier.
In order to emplace musical elements in their corresponding positions, MTN uses two magnitudes. Firstly, for every token a tuple of two integers denotes the staff the element belongs to and its position within the staff. The position is denoted counting the number of steps from the first ledger line below a staff. For those elements without a specific position (such as rests or stems), a null value is used. Secondly, for any object immediately below the class measure, an exact timing value is provided. It is measured in fractions of a quarter note from the start of the measure itself. This information is also provided for every chord in a note group even if this information can be inferred for the sake of simplifying evaluation procedures.
@@
This musical information is necessary for two reasons. The only visual cue for aligning elements in a polyphonic setup is that, at some point, addicional voices have at least one object that is perfectly aligned with one in the main voice. This is very hard to model cleanly with anything other than a graph. Moreover, the main voice of a measure may arbitrarily jump from staff to staff in a system, making timing generally impossible to resolve on a per-staff basis, as can be seen in Figure 3. A small trick to optimise sequence-based notation systems is that one only needs timing information for the first element in a group, as all successive chords’ timings can be inferred from the note types and tuplets attached to them (or their surrounding elements).
Doubly stemmed notes (such as in the last measure of the Arabesques example in Figure 3) can be interpreted like two notes playing at the same time. Therefore, they can be modelled using two distinct groups without needing to break the tree-like structure.
Finally, to enable unambiguous comparison, a reading order of sorts must be established. We propose the following ordering criterion:
@@ 4.3 Extensions
The main goal for the notation is to address symbolic-level OMR. Nevertheless, many practitioners use pipelined architectures where some partial results need to be annotated (most notably, object bounding boxes). Providing these annotations with MTN is straightforward: it is a matter of adding an “annotations” field with each token identifier linked to the desired information. This also includes some other relevant semantic cues such as voices, which we deem unnecessary for the task of score reconstruction but might be useful for some other music information retrieval tasks. A concern surrounding voice information is that its definition around music engraving software is more akin to a layer in a graphic design suite; it is used more as a hacky way of circumventing engraving limitations, making voice information in files like MusicXML rather arbitrary.
@@
The format itself is still maturing phase and open to feedback from the community. We are open to extending it to accommodate the needs of OMR practitioners as they put it into test.
@@ 5 Evaluation Metrics
We propose a set of evaluation metrics that both acknowledge the existence of multiple paradigms for OMR while also setting ways to compare any structured output equally. These metrics draw inspiration from the currently used Symbol Error Rate and ideas from Hajič jr. Hajič jr. et al (2016). We have divided our proposed metrics as tiers depending on the abstraction level they address.
@@ 5.1 Tier 0: Methodology-Specific Metrics
5.2 Tier 1: Primitive Level
The first set of metrics addresses the presence or absence of terminals within the MTN string. The main goal is determining the capacity of models of detecting objects and analysing token co-ocurrence phenomena. These metrics do not take into account structural matters, making them quick to compute. Given a set of unique terminals $T$, which is a subset of the full MTN vocabulary $V$, the multiset of predicted terminals within a sample $P=\left\{p_{1}\dots p_{n}:p_{i}\in T\right\}$ as the multiset of all terminals produced by a model and the multiset of ground truth terminals within a sample $G=\left\{g_{1}\dots g_{m}:g_{i}\in T\right\}$ as the multiset of all terminals existing in the groundtruth (both in no particular order), we define the following metrics:
@@ 5.2 Tier 1: Primitive Level
Since the underlying datasets are quite unbalanced due to the presence of many dominating tokens (e.g. noteheads), these metrics are computed per-class for the entire dataset. In order to produce a single precision and recall measure, results are aggregated per-class using a weighted average, where the weights are the relative frequency of each token in the ground truth.
5.3 Tier 2: Structural Level
This level takes into account the structure of the produced MTN and compares it directly with that of the ground truth. A matching from ground truth elements to those present in the prediction is performed using a tree edit distance algorithm. In particular, since there is a restriction on the ordering of sibling labels, the $O\left(n^{3}\right)$ solution from Zhang and Sasha can be employed Zhang and Shasha (1989).
@@ 5.3 Tier 2: Structural Level
5.4 Tier 3: Semantic Level
This level considers some aspects of music semantics, particularly those concerned with pitch and time. This tier depends on the matching extracted from the structural level in order to identify the association between objects and their ground truth correspondences. Since evaluation has to be possible at the measure level, we use a graphical definition of pitch, which is the position of the notehead within the system.

### v2

@@ 3.4 Evaluation of OMR Models
Evaluating OMR models fairly is important in order to be able to understand the strengths and weaknesses of each system against the rest and being able to dissect the effect of design decisions on the final output.
Unfortunately, most metrics used by the OMR community in modern day are fairly domain-specific and rarely take into account a final representation of music Mengarelli et al (2019). Authors that address OMR as an object detection problem report object detection metrics, whereas authors that address OMR as an image-to-sequence task rely on sequence edit distances. Consequently, there is no realistic way of comparing them since the semantics of the mistakes of the system are lost and the underlying notations are different. The few authors that do report musically-aware metrics restrict themselves to per-note pitch and/or time accuracy Baró et al (2019); Huang et al (2019).
In this direction, there is also debate regarding the specific aspects of music that should be measured. Hajič jr. Hajič jr. (2018) argues that the best metric is one that summarises the entire notation system as one single value, reminiscent of the Symbol Error Rate (SER) metric for strings. We believe that, while such a metric is a good indicator of overall performance (and is definitely useful for benchmarking), it must not be considered on its own, as definite conclusions about the underlying models are very hard to make. Our belief is that there should be an array of metrics that are oriented at different aspects of the score that should be logged together, increasing the coverage of the system. Both cases only make sense under the assumption that comparisons are made on a standardised set of objects and primitives.
@@ 3.5 The Ideal Scope for Recognition
Presently, there is no pre-defined scope to tackle the problem of OMR. Therefore researchers can choose whether to implement recognition at measure Baró et al (2020, 2022), line Calvo-Zaragoza and Rizo (2018a); Baró et al (2019); Wen and Zhu (2022) or page level Pacha et al (2018); Ríos-Vila et al (2022a); Zhang et al (2023). While there are practical advantages on each perspective, there is an arguing point that should be brought to discussion given the current situation of OMR and Computer Vision as a whole.
Current state-of-the-art object detection and recognition models struggle with both small Yu et al (2020); Pacha (2021); Baró et al (2022) and scale-variant symbols Zhang et al (2018). Music scores are quintessential examples of both phenomena: large symbols (slurs, ties, staves) coexist with very small ones (dots, fingering numbers), and writers draw symbols at their preferred relative scale without affecting their meaning – something that is particularly true for noteheads and stems – while also writing semantically-distinct symbols that change based on their size – grace notes. Additionally, the number of objects present in the input image also affects the recognition performance, particularly when objects accumulate in a limited region of the image.
Considering this, it should be possible to tackle OMR from the smallest scope possible: the measure level. Addressing higher scopes can be achieved by composing smaller semantic units. The downside is that this requires more effort to be put in the analysis of the layout of the scores, as all successive recognition steps depend on finding the measures reliably on the page. Nevertheless, it can be argued that it is an objectively simpler problem than recognition: staves and bar lines are very prominent, very ubiquitous and can easily be identified heuristically Fujinaga (2004); Cardoso et al (2009); Fornés et al (2013); Bui et al (2014); Calvo-Zaragoza et al (2017a); Egozy and Clester (2022), as opposed to many of the other musical symbols.
@@ 4 The Music Tree Notation Format
  • simplifies conversion to a final structured format,
  • enables comparison of diverse OMR methods on equal grounds and
@@ 4.1 Rationale
The basis to our proposal for an OMR framework for inference, score reconstruction and evaluation is the Music Tree Notation format (MTN). We believe the final goal of structured OMR is the reconstruction of the score at the visual domain. The core idea of this format is therefore to build a notation that exclusively models relationships between graphical symbols and defers inference of music semantics until a later stage. Only those high-level music concepts that are strictly required to reconstruct the score unambiguously are kept if and only if there is a direct graphical cue that allows straightforward inference.
Score reconstruction should be possible regardless of the scope of recognition of the underlying system – page, line or measure. Nevertheless, a minimal semantic unit is still required in order to reconstruct music semantics afterwards and facilitate alignment of predictions and their groundtruth. Exploiting the assumption OMR is performed on the visual domain and therefore application of semantic-altering objects is irrelevant, we propose using the measure as the minimal unit of music on which to evaluate. Moreover, reconstruction should be possible regardless of the origin of the transcription. The format must not require localisation of symbols – only their relationships – in order to accommodate the use of end-to-end or pipelined techniques interchangeably.
MTN should act as a final target for OMR, through which to convert the score into any other high-level notation system. Employing a tree structure simplifies developing converters to new formats, which can be implemented by writing a tree traversal algorithm at linear complexity. This has the potential of facilitating OMR systems from research being deployed into real use-cases as well as enabling any source of structured data to be used in OMR applications. Therefore, MTN draws heavy inspiration from both MEI and MusicXML, but “normalises” their structure in such a way that every score can only be expressed in one way. This also allows implementation many of the already-studied evaluation requirement ideas from other authors, particularly Hajič jr. Hajič jr. et al (2016) and Byrd Byrd and Simonsen (2015).
Finally, music notation is very rich and tends to bend its own rules quite heavily. The goal for this notation and evaluation system is to accommodate a most significant subset of CWMN, but covering every single exception proves to be a daunting task. Byrd Byrd and Simonsen (2015) presents some specific cases such as a spanning slur with seven inflection points that is rather complicated to model in the logical domain. Therefore, our rationale is that support for these should not be a priority  [^5] This specific change could potentially be modelled in MTN using the tools already in the format either by providing the bounding box of the slur (which would be rather imprecise) or its pixel mask associated to its ID in the tree.. We believe it is better to have some solid tools on a core subset of music rather than trying to cover every possible case with highly complex and fragile ones as this makes adoption of the format much more difficult. In this regard, the format is designed so it is very easy to add support for new types of music primitives or add new sources of information.
@@ 4.2 Description of the Format
A graphical representation of a simple measure engraved in MTN can be seen in Figure 2. The core element of this format is the Musical Primitive, a concept that is quite widespread in the OMR literature Baró et al (2016, 2020); Calvo-Zaragoza and Rizo (2018a); Tuggener et al (2018) and can be defined as any of the independent structural elements that may or may not be combined together to form a semantic unit in the music score. The set of musical primitives includes all graphical elements in a score that are self-contained and require no other symbols to convey meaning (this includes rests, clefs or time signature symbols), the set of graphical elements that compose notes (noteheads, stems, flags, dots, accidentals, etc.) and other miscellaneous elements such as numbers for compound time signatures. Every primitive is given a unique work-level identifier.
These primitives associate together to form more abstract constructs. This is modelled in MTN using a tree-like structure of higher-order elements resembling an Abstract Syntax Tree (AST), which defines the set of dependencies among objects in the score. This idea emulates parsing the contents of the score using a grammar, enabling the bulk of tools and research on parsers, parser generators and AST analysis and processing to be used in the context of music. Furthermore, it is a structure that can be modelled very easily using an exchange format such as XML. The structure of the format may be found in Table 2.
Tree-based music representations for evaluation of recognition has been explored before in Foscarin et al (2019). In this paper they also nest tuplets within the hierarchy. Nevertheless, this is not possible for any tuplet in MTN because they can extend beyond the scope of the current group and incorporate neighboring rests or other objects, which would break the tree assumption of the format. Instead, this is modelled using two start/end tokens.
Table 2: Definition of the main abstract elements within MTN and the overall structure of the format. Each of the lowest levels of the notation contain the various final tokens. For more information, refer to the implementation of the format.
@@
Attributes Any semantic change the score goes through.
Staff Group of changes occuring on the same staff.
Clef A change of clef in the score.
TimeSig A change of time signature in the score.
Key A change of key in the score.
Barline A measure separator and its attached modifiers.
Direction A single playback directive.
Note Group Group of notes tied together by at least one beam.
Note Group Note Groups can be nested, each owning one or more beams
Chord Set of notes in unison attached to the same stem, if it exists.
Stem Stem direction information and beams or flags attached to it.
Note A notehead and its attached modifiers.
Rest Rest token and its attached modifiers.
There are some elements in music that break the tree-like structure assumption. These are elements that connect multiple notes together outside their local note group structure: slurs, ties, parentheses and tuplets, among others. Both MEI and MusicXML acknowledge this limitation and circumvent it through the use of identifiers. MTN is no different: it provides a unique starting and ending token for each side of the object and gives both ends the same identifier.
In order to describe the position of MTN elements, two magnitudes are used. Firstly, for every token a tuple of two integers denotes the staff the element belongs to and its position within the staff. The position is denoted counting the number of steps from the first ledger line below a staff. For those elements without a specific position (such as rests or stems), a null value is used. Secondly, for any object immediately below the class measure, an exact timing value is provided. It is measured in fractions of a quarter note from the start of the measure itself. This information is also provided for every chord in a note group even if this information can be inferred for the sake of simplifying evaluation procedures.
@@
This musical information is necessary for two reasons. The only visual cue for aligning elements in a polyphonic setup is that, at some point, additional voices have at least one object that is perfectly aligned with one in the main voice. This is very hard to model cleanly with anything other than a graph. Moreover, the main voice of a measure may arbitrarily jump from staff to staff in a system, making timing generally impossible to resolve on a per-staff basis, as can be seen in Figure 3. A small optimisation for sequence-based notation systems is that only the first element in a group needs timing information, as all successive chords’ timings can be inferred from the note types and tuplets attached to them (or their surrounding elements).
Doubly stemmed notes (such as in the last measure of the Arabesques example in Figure 3) can be interpreted as two notes playing at the same time. Therefore, they can be modelled using two distinct groups without needing to break the tree-like structure.
Finally, to produce unambiguous scores, a reading order of sorts must be established. We propose the following ordering criterion:
@@ 4.3 Extensions
The main goal for the notation is to address symbolic-level OMR. Nevertheless, many practitioners use pipelined architectures where some partial results need to be annotated (most notably, object bounding boxes). Providing these annotations with MTN is straightforward: it is a matter of adding an “annotations” field with each token identifier linked to the desired information. This also includes some other relevant semantic cues such as voices, which we deem unnecessary for the task of score reconstruction but might be useful for some other music information retrieval tasks. A concern surrounding voice information is that its definition around music engraving software is more akin to a layer in a graphic design suite; it is mostly used as a way of circumventing engraving limitations, making voice information in files like MusicXML rather arbitrary.
@@
The specification of the format provides ways of developing new extensions as new needs arise.
@@ 5 Evaluation Metrics
We propose a set of evaluation metrics that both acknowledge the existence of multiple paradigms for OMR while also setting ways to compare any structured output equally. These metrics draw inspiration from the currently used Symbol Error Rate and some ideas from Hajič jr. Hajič jr. et al (2016). We have divided our proposed metrics as tiers depending on the abstraction level they address and the problems they can help diagnose.
@@ 5.1 Tier 0: Methodology-Specific Metrics
5.2 Tier 1: Primitive detection
The first set of metrics addresses the presence or absence of terminals within the MTN string. The main goal is determining the capacity of models of detecting objects and analysing token co-ocurrence phenomena. This is especially important in OMR because the distribution of classes in music is extremely unbalanced: Noteheads are orders of magnitude more common than almost any other token. Understanding what tokens are missed and why is key to developing a good recognition system, since missing a single stem can alter the reconstruction of the score completely.
These metrics do not take into account structural matters, making them quick to compute. Given a set of unique terminals $T$, which is a subset of the full MTN vocabulary $V$, the multiset of predicted terminals within a sample $P=\left\{p_{1}\dots p_{n}:p_{i}\in T\right\}$ as the multiset of all terminals produced by a model and the multiset of ground truth terminals within a sample $G=\left\{g_{1}\dots g_{m}:g_{i}\in T\right\}$ as the multiset of all terminals existing in the groundtruth (both in no particular order), we define the following metrics:
@@ 5.2 Tier 1: Primitive detection
These metrics are computed per-class for the entire dataset. In order to produce a single precision and recall measure, results are aggregated per-class using a weighted average, where the weights are the relative frequency of each token in the ground truth.
5.3 Tier 2: Structure Reconstruction
This tier takes into account the structure of the produced MTN and compares it directly with that of the ground truth. A matching from ground truth elements to those present in the prediction is performed using a tree edit distance algorithm. In particular, since there is a restriction on the ordering of sibling labels, the $O\left(n^{3}\right)$ solution from Zhang and Sasha can be employed Zhang and Shasha (1989). In practice, we use a Python implementation Joa (2017) of Pawlik et al.’s APTED algorithm Pawlik and Augsten (2016).
@@ 5.3 Tier 2: Structure Reconstruction
5.4 Tier 3: Semantic Reconstruction
This tier considers whether the subset of music semantics required by MTN has been extracted correctly. It depends on the matching extracted from the structural level in order to identify the association between objects and their ground truth correspondences. Since evaluation has to be possible at the measure level, the graphical definition of pitch is used, which is the position of the notehead within the system.